# Tunable lower critical fractal dimension for a non-equilibrium phase transition

### v1

@@ Abstract
We theoretically investigate the role of spatial dimension and driving frequency in a non-equilibrium phase transition of a driven-dissipative interacting bosonic system. In this setting, spatial dimension is dictated by the shape of the external driving field. We consider both homogeneous driving configurations, which correspond to standard integer-dimensional systems, and fractal driving patterns, which give rise to a non-integer Hausdorff dimension for the spatial density. The onset of criticality is characterized by critical slowing down in the excited density dynamics as the system asymptotically approaches the steady state, signaling the closing of the Liouvillian frequency gap. We show that the lower critical dimension —below which the non-equilibrium phase transition ceases to occur— can be controlled continuously by the frequency detuning of the driving field.
Introduction — A phase transition [1] for a physical system manifests as a discontinuous behavior of observable quantities emerging in some thermodynamic limit. In this context, the spatial dimension plays a key role in determining critical behavior. A particularly important concept in this regard is the lower critical dimension $d_{L}$: a phase transition that occurs for dimension $d\geq d_{L}$ ceases to exist for $d<d_{L}$, as fluctuations in lower-dimensional systems tend to suppress singularities.
Non-equilibrium [2] phase transitions are critical phenomena that remain less explored and understood compared to their equilibrium counterparts . A critical point of a non-equilibrium system can be characterized by a critical slowing down of the system dynamics approaching the steady-state in a thermodynamical limit [3]. In the Liouvillian framework, this is equivalent to studying the closing of the frequency gap associated to the master equation describing the non-equilibrium dynamics of the density matrix [3]. Interacting optical systems [4] are particularly well-suited for studying such transitions due to the precise control of external driving, nonlinearities and the accurate characterization of losses, which compete in shaping non-trivial non-equilibrium dynamics and steady-state [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. For interacting bosons such as photons, in addition to the usual thermodynamic limit of large spatial size, critical behavior emerges in another thermodynamic limit when the physical parameters are tuned in such a way that the number of bosons in the state becomes arbitrarily large. This has been shown to occur both theoretically and experimentally in driven systems consisting of a single two-level atom coupled to single cavity photon mode [6, 8] or a driven single cavity mode with Kerr nonlinearity [15, 16, 17], achieved in the limit of weak photon nonlinearity. Here, we will focus primarily on the conventional thermodynamic limit of large spatial size, which, in the context of optics, corresponds to an infinite number of modes.
While the role of integer spatial dimensions in equilibrium critical phenomena is well studied, the influence of non-integer dimensions, particularly in fractal patterns, is not as much understood even in equilibrium systems [18, 19, 20, 21]. Notably, studies on the equilibrium Ising model [22] suggested that the critical behavior is not solely determined by the fractal (Hausdorff) dimension but also other geometrical attributes [22, 23, 24, 25].
@@
Recent studies on driven-dissipative interacting bosonic systems have revealed critical behavior in the limit of large spatial size, particularly in two-dimensional (2D) systems [26, 27, 28]. In these systems, a first-order phase transition manifests as a discontinuous jump in the bosonic density at a critical driving intensity. A theoretical study demonstrated that, within a range of frequency detunings close to the dissipation rate, these systems exhibit an emergent equilibrium behavior that can be mapped onto the equilibrium Ising model [26]. In this analogy, the high- and low-density phases correspond to the two states of the Ising variable. Since the Ising model is known to undergo a phase transition in 2D but not in 1D, a similar dimensional dependence is expected for these driven-dissipative systems. A subsequent study on continuous systems, where the effective dimension is imprinted by a properly shaped optical driving field [28], confirmed this behavior: a phase transition emerges in 2D but remains absent in 1D when the frequency detuning is close to the linear-regime dissipation rate. However, the role of frequency detuning was not explored further. Additionally it remained unclear what behavior would characterize the system at intermediate dimensions, that in continuous systems can be naturally obtained by shaping the driving field. Indeed, the impact of fluctuations due to many-mode dynamics strongly depends on the spatial dimension.
Several fundamental questions have never been considered: What happens to this class of non-equilibrium phase transitions in fractional dimensions? Does the system exhibit a lower critical dimension below which no phase transition occurs? How does frequency detuning influence this critical dimension? Addressing these fundamental questions is central to the present work, as they may unveil new paradigms in non-equilibrium critical behavior.
In this Letter, we investigate the phase transition of a driven-dissipative, continuous, and interacting 2D bosonic gas. We use both homogeneous and fractal-shaped driving patterns to impose different Hausdorff dimensions. We demonstrate that critical behavior is strikingly governed by both the Hausdorff dimension and the frequency detuning of the driving.
@@
$i\frac{d\psi}{dt}=\left[-\Delta-\frac{\hbar\nabla^{2}}{2m}+g\left|\psi\right|^{2}-i\frac{\gamma}{2}\right]\psi+iF(\textbf{r})\,.$ (1)
@@
Here, $\psi(\textbf{r},t)$ denotes the mean value of the bosonic field operator, depending on the two-dimensional spatial position $\mathbf{r}$ and time $t$. The equation is in the so-called pump rotating frame, where the frequency detuning $\Delta$ represents the difference between the driving frequency and the frequency of the boson mode with zero wavevector (corresponding to the lowest frequency mode). The nonlinear repulsive interaction is quantified by $g$, while $\gamma$ characterizes the natural dissipation rate of the system. In the following, we will express lengths in units of the characteristic length $\xi=\sqrt{\frac{\hbar}{2m\gamma}}$. Finally, $F(\textbf{r})$ denotes the spatially dependent driving field.
@@
In this class of systems, a phase transition manifests in a sharp increase of the bosonic density $n=|\psi|^{2}$ when the driving intensity $I=|F|^{2}$ reaches a critical value $I_{c}$ [26, 27, 28]. This is related to the bistable behaviour in the mean-field treatment for detunings $\Delta>\sqrt{3}/2~{}\gamma$ [4, 15], whose robustness to fluctuations depends on dimension. Above the critical dimension, a sharp jump of the density is obtained in the thermodynamic limit of an infinite system size. This corresponds to a first-order phase transition from a dark (low boson density) to a bright (high boson density) state.
@@
The effective spatial dimensional can be controlled with the spatial shape of the driving field $F(r)$: pumping uniformly over a region of the plane results in a two-dimensional system with diffusive boundary conditions (see first column of Fig. 1), while imposing a different geometry results in a lower dimension. For example, a pump profile that is narrow in one direction and elongated in the other creates an effectively one-dimensional system [28] (fourth column of Fig. 1). But this can be extended to non-integer dimensions if the driving field has a fractal shape, as exemplified in the second and third columns of Fig. 1, where a Sierpinski carpet and a Sierpinski cross with a finite number of iterations are shown.
@@
Note that the resulting bosonic density has in general a slightly distorted shape, as shown in the lower panels of Fig. 1 (more patterns are shown in the Supplementary Material [30]). For example, in the second column, smaller holes in the Sierpinski carpet are partially filled and more affected than larger holes. This is due mainly to interference and dispersive effects, as well as the nonlinear boson-boson interactions, that introduce intrinsic length scales in the system. Despite these distortions and the finite size, the resulting patterns possess their own fractal dimension (slightly different from the nominal one of the ideal, infinitely self-similar patterns) that can be quantified with the box-counting method [31, 30] (for more details see  [^1] This method quantifies the fractal dimension by overlaying a grid of $s^{2}$ square boxes, each of edge length $L/s$, where $L$ is the system size. After applying a thresholding algorithm [34] to extract the structure from a grayscale image, the number of boxes $A$ that intersect the pattern is counted. Assuming a scaling relation $A\sim s^{d_{H}}$, the fractal (Hausdorff) dimension $d_{H}$ is obtained as the slope of $\log A$ versus $\log s$ over a range of scales. More details are provided in the Supplementary Materials [30].). In Fig. 1, we have annotated both the pump pattern and the resulting densities with their respective box-counting dimensions $d_{H}$. Importantly, even in the presence of distortions and finite-size effects, the extracted dimensions remain fractional and close to the nominal one.
@@
While for integer-dimensional configuration the system size can be changed continuously, reaching the thermodynamic limit of a fractal system is less straightforward and can be approached in several ways. Here, we consider the limit obtained by increasing the total system size while keeping the length of the smallest structural detail $l$ constant, as illustrated in Fig. 2(a). In all calculations reported in this work, we take $l=3.45\xi$. If this length were allowed to vary, the smallest elements of the pattern would change with system size, thereby changing its ratio to the intrinsic physical length scales and drastically altering its box counting dimension. Moreover, to maintain the fractal structure as the system grows, at each step we also increase the number of self-similar iterations. For example for the Sierpinski carpet, this implies that the linear size triples with each additional iteration, as depicted in Fig. 2(a). Although this exponential growth poses computational challenges, we were able to reach sufficiently large system sizes to observe the emergence of the asymptotic behavior characteristic of the thermodynamic limit.
@@
Phase transition markers — In the investigation of phase transitions, several markers serve as key indicators for identifying their presence or absence. One such marker is the maximal slope of the spatially averaged steady-state density as a function of the input intensity. In particular, when a first-order phase transition occurs, this slope diverges, signaling a discontinuous change in the system’s behavior [28]. Another crucial indicator is the asymptotic decay rate $\lambda$, also referred to as the Liouvillian gap in the Lindbladian framework [5, 3]. At long times, the system relaxes exponentially towards its steady state with a decay rate $\lambda$, whose inverse is the characteristic timescale for equilibration. At the critical point of a driven-dissipative phase transition, $I=I_{c}$, one expects critical slowing down, namely the system requires an infinite amount of time to reach its steady state. In other words, at critical driving the asymptotic decay rate $\lambda$ tends to zero in the thermodynamic limit [3, 15]. Conversely, if in this infinite-size limit the decay rate remains finite for all values of the driving intensity, no phase transition occurs.
@@
The asymptotic decay rate $\lambda$ can be obtained from the time-dependence of the spatially averaged density $\bar{n}(t)=\int\mathrm{d}^{2}x\ n(t,x)$, that at long times will evolve as ${\bar{n}(t)-\bar{n}_{ss}\propto e^{-\lambda t}}$, where $\bar{n}_{ss}$ is its steady-state value. We compute this value by numerically evolving the dynamical equation (1) and performing a Fourier analysis of $\bar{n}(t)$. This procedure is repeated for different driving intensities $I$, to find the smallest value of $\lambda$, i.e. the minimal asymptotic decay rate $\lambda_{\min}=\min_{I}\lambda(I)$. To probe the presence of a critical behaviour, we compare the values of $\lambda_{\min}$ obtained for different system sizes: a critical point is present if this quantity tends to zero for increasing system sizes, while it is absent if if saturates to a finite value.
@@
Figure 2: (a) Illustration of the approach to the thermodynamic limit by expanding the size of a fractal pattern. At each step, the total linear size $L$ is multiplied by three, while the smallest resolved feature $l$ remains constant. (b–d): Minimal asymptotic decay rate $\lambda_{\min}$ as a function of the linear system size $L$ for detuning values $\Delta=1.0\gamma$, $\Delta=2.0\gamma$, and $\Delta=3.0\gamma$. Different markers correspond to different spatial geometries indicated in the legend.
@@
Results at fixed detuning — We begin by examining the well-studied detuning $\Delta=\gamma$, for which in previous investigations the first-order phase transition was found to occur in a two-dimensional system, and to be absent in a one-dimensional one [26, 27, 28]. This is confirmed by our calculations of $\lambda_{\min}$, whose dependence on the system size is shown in Fig. 2(b), with green squares for a 2D system and orange bars for a 1D system. Here and later, for each geometry we computed $\lambda_{\min}$ for a range of driving intensities and identified the minimal value. In fact, for the 2D case and in the limit of large systems, $\lambda_{\min}$ vanishes (with a precision given by the numerical resolution, which is set by the final time of the calculated temporal dynamics). For the considered detuning, this critical slowing down is instead absent in the one-dimensional system, as $\lambda_{\min}$ saturates to a finite value.
@@
We extended this investigation to fractal dimensions. For example, for the Sierpinski carpet (second column of Fig. 1), $\lambda_{\min}$ has a minimum around the expected mean-field critical driving [30], but the value of this minimum converges to a finite value for increasing system size, as shown in Fig. 2(b) with empty blue square markers. The system therefore exhibits slowing down, but it is not critical. The shape of this slowest mode is non-trivial and will be the object of further studies [30]. However we can say that, for this value of the detuning, the Sierpinski carpet of dimension $d_{H}\simeq 1.84$ is below the critical dimension. This can be also characterized in terms of the steady-state average density, that does not display a discontinuous jump as a function of $I$, and has instead a smooth crossover from the low density to the high density phase [30]. We obtain a similar behavior for the Sierpinski cross (third column of Fig. 1), which has $d_{H}\simeq 1.5$ and whose convergence of $\lambda_{\min}$ to a finite value is shown with red cross markers in Fig. 2(b).
@@
Figure 3: Minimal asymptotic decay rate $\lambda_{\min}$ versus Hausdorff dimension for a fixed detuning $\Delta=\gamma$. The driving patterns associated with the various points are visualized in the Supplementary Material. Horizontal error bars show the uncertainty of the box-counting method. System sizes are reported in the main text.
@@
The actual value of $\lambda_{\text{min}}$ also quantifies how far the system is from criticality, and we found it to depend strongly on the system’s dimensionality. This is shown in Fig. 3, where we report the minimal asymptotic decay rate as a function of the Hausdorff dimension $d_{H}$, for a variety of fractal geometries (the patterns for the light blue circular markers are shown in [30]). These values are the ones for the largest computationally feasible sizes (ranging from $L=81l$ to $L=216l$, depending on the geometry), that are representative of the thermodynamic limit $\lim_{L\to\infty}\lambda_{\text{min}}$, since we assured convergence of this value. The difference between $\lambda_{\min}$ for the two largest considered sizes gives us a vertical error bar that is always smaller than the marker size. The horizontal error bars are instead the moderate uncertainty in the Hausdorff dimension given by the box-counting method, which is more significant for patterns that have a fractal dimension smaller than one (Cantor sets). Remarkably, we have found that $\lambda_{\min}$ exhibits a clear monotonous dependence on the Hausdorff dimension $d_{H}$, with the slowing down continuously approaching a critical one while the dimension is increased. In this case $\Delta=\gamma$, we observe that the decay rate only tends to zero when the dimension tends to two, i.e. the lower critical dimension for the first-order phase transition is $d_{L}(\Delta=\gamma)=2$. This extends previous results [26, 27, 28] to fractional dimensions, but, as we will show in the following, we have made the remarkable discovery that the lower critical dimension surprisingly decreases significantly from 2 as the detuning $\Delta$ increases.
@@
Figure 4: Minimal asymptotic decay rate as a function of detuning $\Delta$ for different spatial dimensions. All data points are obtained for a system size $L=81l$. The vertical error bars represent the difference in $\lambda_{\min}$ between this size and $L=27l$, providing an estimate of the finite-size effects and our proximity to the thermodynamic limit. Inset: Critical detuning $\Delta_{c}$ as a function of the Hausdorff dimension $d_{H}$. Error bars correspond to the numerical resolution of our simulation.
@@
Changing the lower critical dimension — As shown in Figs. 2(c)-(d) and in Fig. 4, increasing the driving detuning $\Delta$ leads to a systematic decrease of $\lambda_{\min}$, and eventually to its vanishing, for all geometries. This implies that the system is critical and exhibits a first-order phase transition even for dimensions $d_{H}<2$, provided that the ration $\Delta/\gamma$ is large enough. We define the critical detuning $\Delta_{c}$ as the value above which $\lambda_{\min}$ vanishes in the thermodynamical limit. This value is geometry-dependent: for the 2D system (green squares), it is $\Delta_{c}\simeq 1.0\gamma$, close to the known previously studied case. The critical detuning $\Delta_{c}$ is larger for lower-dimensional systems. This can be seen in Fig. 2(c), where, for large enough systems, $\lambda_{\min}$ vanishes not only for the 2D square, but also for the Sierpinski carpet, which for $\Delta=\gamma$ is not critical. Correspondingly, the smooth crossover in the average density for $\Delta=\gamma$, becomes a sharp discontinuity [30]. For $\Delta=2\gamma$ the lower-dimensional Sierpinski cross and the one-dimensional system are still non-critical, but both become for a higher $\Delta=3\gamma$, as shown in Fig. 2(d). These results show that it is possible to tune the lower critical dimension, or, in other words, to make the phase transition occur at any fixed dimension by increasing $\Delta$. The inset of Fig. 4, shows that the lower critical dimension decreases continuously as the Hausdorff dimension is changed. Interestingly, also the 1D case and fractals with dimension lower than one (e.g., Cantor sets) can be made critical by increasing $\Delta/\gamma$ enough. In fact, in the limit $d_{H}\to 0^{+}$, we can expect $\Delta_{c}\to+\infty$. Indeed, for a zero-dimensional (one-mode) nonlinear bosonic resonator, it is known that the asymptotic decay rate vanishes only in the limit of infinite detuning [16, 33].

### v2

@@ Abstract
We theoretically investigate the role of spatial dimension and driving frequency in a non-equilibrium phase transition of a driven-dissipative interacting bosonic system. In this setting, spatial dimension is dictated by the shape of the external driving field. We consider both homogeneous driving configurations, which correspond to standard integer-dimensional systems, and fractal driving patterns, which give rise to a non-integer Hausdorff dimension for the spatial density. The onset of criticality is characterized by critical slowing down in the excited density dynamics as the system asymptotically approaches the steady state. By analyzing the system-size dependence of the asymptotic decay rate using numerical simulations of the full multi-mode dynamics, complemented by an analytical statistical mean-field treatment, we determine the lower critical dimension of the non-equilibrium phase transition. We show that this dimension can be non-integer and fractal in nature, and that it can be tuned continuously via the frequency detuning of the driving field.
Introduction — A phase transition [21] in a physical system manifests as a nonanalytic or discontinuous behavior of observable quantities that emerges in an appropriate thermodynamic limit. Spatial dimensionality plays a central role in determining critical behavior, with a particularly important concept being the lower critical dimension $d_{L}$: a phase transition that exists for dimensions $d\geq d_{L}$ is suppressed by fluctuations and ceases to occur for $d<d_{L}$. While equilibrium critical phenomena are well understood when restricted to integer spatial dimensions, the situation is far less settled for non-integer dimensions, especially in systems defined on fractal geometries [22, 17, 34, 24]. In particular, studies of the equilibrium Ising model [16] have shown that critical behavior is not determined solely by the fractal (Hausdorff) dimension, but can also depend sensitively on additional geometrical characteristics of the underlying structure [16, 7, 26, 33].
Further complications arise out of equilibrium, where phase transitions are substantially less explored and understood than their equilibrium counterparts. In non-equilibrium systems, a critical point is characterized by a critical slowing down of the dynamics as the steady state is approached in the thermodynamic limit. For open quantum systems [5], this behavior is directly associated with the closing of the spectral gap of the Liouvillian governing the system dynamics [23]. Interacting optical systems [8] provide an especially powerful platform for investigating such transitions, owing to the precise control of external driving, nonlinearities, and dissipation. These ingredients compete to generate rich non-equilibrium dynamics and steady states [18, 6, 2, 12, 3, 4, 14, 31, 28, 32]. While for interacting bosonic systems such as photons the thermodynamic limit may also be achieved by taking the occupation number of a mode to be arbitrarily large [6, 12, 9, 27, 13], here we instead focus on the conventional thermodynamic limit of large spatial extent, which in optical settings corresponds to an infinite number of modes.
Recent studies of driven-dissipative interacting bosonic systems have revealed a first-order phase transition in two dimensions (2D) [15, 30, 20], manifested by a discontinuous jump in the bosonic density. A theoretical analysis considered the regime in which the frequency detuning of the drive from the bosonic resonance is comparable to the dissipation rate and demonstrated that the system can display an emergent equilibrium behavior that maps onto the equilibrium Ising model [15]. In this correspondence, the low and high density phases play the role of the paramagnetic and ferromagnetic phases, respectively. Since the Ising model exhibits a phase transition in 2D but not in 1D, an analogous dimensional dependence would be naturally expected for these driven dissipative systems. This expectation was subsequently experimentally confirmed on continuous platforms, where the detuning was set equal to the dissipation rate and the effective spatial dimension was engineered through a suitably shaped optical driving field, realizing one- and two-dimensional systems with diffusive boundary conditions [20].
Despite these advances, fundamental questions concerning non-equilibrium critical phenomena in driven-dissipative systems remain open. In particular, the general role of spatial dimensionality is far from established, as existing studies have explored only a restricted region of parameter space. Crucially, previous investigations have focused on the specific regime in which the frequency detuning is close to the dissipation rate, leaving open the question of whether the observed dimensional dependence persists away from this fine-tuned condition. At the same time, continuous driven systems offer a unique degree of control over spatial structure, as the effective dimensionality can be engineered directly through the spatial profile of the driving field. This makes it possible to go beyond integer dimensions and to systematically investigate critical behavior in fractional, non-integer spatial dimensions, a regime that has so far remained unexplored.
@@
In this Letter, we show that in a non-equilibrium setting the lower critical dimension is a tunable quantity and can take fractional values. We establish this by studying the phase transition of a driven dissipative, continuous, interacting two dimensional bosonic gas subjected to either homogeneous or fractal shaped driving fields, which impose different Hausdorff dimensions on the system. By analyzing the system size dependence of the long time relaxation rate toward the steady state, we demonstrate that the existence of a critical point is jointly controlled by the fractal dimension and the frequency detuning of the drive. These findings reveal that non equilibrium phase transitions can exhibit a level of richness and controllability that has no direct counterpart in equilibrium critical phenomena and remains largely unexplored.
@@
$i\frac{d\psi}{dt}=\left[-\Delta-\frac{\hbar\nabla^{2}}{2m}+g\left|\psi\right|^{2}-i\frac{\gamma}{2}\right]\psi+F(\textbf{r})\,.$ (1)
@@
Here, $\psi(\textbf{r},t)$ denotes the mean value of the bosonic field operator, which depends on the spatial position $\mathbf{r}$ and the time $t$, and $n(\textbf{r},t)=|\psi(\textbf{r},t)|^{2}$ is the bosonic density. The equation is written in the so-called pump rotating frame, with $\Delta$ the detuning between the driving frequency and the frequency of the spatially uniform, lowest-frequency bosonic mode. Moreover, $g$ quantifies the nonlinear repulsive interaction, $\gamma$ is the natural dissipation rate of the system, and $F(\textbf{r})$ denotes the spatially dependent driving field. In the following, we express lengths in units of a characteristic length $\xi=\sqrt{\hbar/(2m\gamma)}$. In all simulations we use $g/(\gamma\xi^{2})=0.0431$. Note that, within the non-equilibrium Gross-Pitaevskii framework, changing its value is equivalent to a rescaling of the driving field and bosonic density, hence not affecting the presence or absence of a phase transition [1].
@@
In this class of systems, a phase transition is signaled by a sharp increase of the spatially averaged steady state density $\bar{n}_{ss}$ when the driving intensity $I=|F|^{2}$ reaches a critical value $I_{c}$ [15, 30, 20]. At the level of a mean-field description of a single-mode driven-dissipative bosonic system, no genuine phase transition is possible. Instead, for detunings $\Delta>\sqrt{3}/2\gamma$, the mean-field equations display bistability between low and high-density steady state solutions [8, 9, 11]. This bistable behavior does not correspond to a phase transition but represents a mean-field precursor of phase coexistence. Whether such coexistence survives and criticality emerges in the thermodynamic limit is decided by fluctuations associated with the full multi-mode dynamics. In particular, the criticality depends crucially on the spatial dimension of the system. Above a critical dimension, fluctuations are sufficiently suppressed and a discontinuous jump of the density persists in the thermodynamic limit, signaling a first-order phase transition from a dark low density state to a bright high density state.
@@
The effective spatial dimensional can be controlled with the spatial shape of the driving field $F(r)$: pumping uniformly over a region of the plane results in a two-dimensional system with diffusive boundary conditions (first column of Fig. 1), while imposing a different geometry results in a lower dimension. For example, a pump profile that is narrow in one direction and elongated in the other creates an effectively one-dimensional system [20] (fourth column of Fig. 1). This can be extended to non-integer dimensions if the driving field has a fractal shape, as exemplified in the second and third columns of Fig. 1, where a Sierpinski carpet and a Sierpinski cross with a finite number of iterations are shown.
@@
Note that the resulting bosonic density has in general a slightly distorted shape, as shown in the lower panels of Fig. 1 (more patterns are shown in the Supplementary Material [1]). For example, in the second column, smaller holes in the Sierpinski carpet are partially filled and more affected than larger holes. This is due to interference and dispersive effects, as well as boson-boson interactions, that introduce intrinsic length scales in the system. Despite these distortions and the finite size, the resulting patterns possess their own fractal dimension, that can be quantified with the box-counting method  [^1] The box-counting method [29] quantifies the fractal dimension by overlaying a grid of $s^{2}$ square boxes, each of edge length $L/s$, where $L$ is the system size. After applying a thresholding algorithm [25] to extract the structure from a grayscale image, the number of boxes $A$ that intersect the pattern is counted. Assuming a scaling relation $A\sim s^{d_{H}}$, the fractal (Hausdorff) dimension $d_{H}$ is obtained as the slope of $\log A$ versus $\log s$ over a range of scales. More details are provided in the Supplementary Materials [1]. and remain fractional and close to the one of the ideal infinite self similar pattern. The box counting dimension $d_{H}$ of the driving patterns and the resulting densities are reported in Fig. 1.
@@
While for integer-dimensional configuration the system size can be changed continuously, reaching the thermodynamic limit of a fractal system is less straightforward. Here, we consider the limit obtained by increasing the total system size while keeping the length of the smallest structural detail $l$ constant, as illustrated in Fig. 2(a). This is important to set its ratio to the intrinsic physical lengths, which would otherwise drastically alter the box counting dimension of the resulting bosonic density. In all calculations we take $l=10\mu m=3.45\xi$ as an example such that $l>\xi$. This also implies that, at each step, the number of self-similar iterations is increased by one. This procedure results in a system size that increases by integer factors, as shown for the Sierpinski carpet in Fig. 2(a), for which the linear size triples at each step. Although this exponential growth poses computational challenges, we were able to reach sizes that are much larger than the ones typically implemented in experimental realizations [20], and for which the critical behavior in 2D has already emerged.
@@
Phase transition markers — In the investigation of phase transitions, several markers serve as indicators for identifying their presence or absence. In our case, one could use the maximal slope of the spatially averaged steady-state density as a function of the input intensity, that diverges when a first-order phase transition occurs [20]. Another useful indicator is the asymptotic decay rate $\lambda$, also referred to as the Liouvillian gap in the Lindbladian framework [18, 23], that is the decay rate with which the system relaxes exponentially towards its steady state at long times. At the critical point of a driven-dissipative phase transition, critical slowing down occurs, that is the system requires an infinite amount of time to reach its steady state, and $\lambda$ tends to zero in the thermodynamic limit [23, 9]. Conversely, if in this infinite-size limit the decay rate remains finite for all values of the driving intensity, no phase transition occurs.
@@
The asymptotic decay rate $\lambda$ can be obtained from the time-dependence of the spatially averaged density $\bar{n}(t)=\int\mathrm{d}^{2}x\ n(t,x)$, that at long times evolves as $\bar{n}(t)-\bar{n}_{ss}\propto e^{-\lambda t}$}, where $\bar{n}_{ss}$ is the steady-state value. We compute this value by numerically evolving the dynamical equation (1) and performing a Fourier analysis of $\bar{n}(t)$ [1]. This procedure is repeated for different driving intensities $I$ to find the minimal asymptotic decay rate $\lambda_{\min}=\min_{I}\lambda(I)$. To probe the presence of a critical behaviour, we compare the values of $\lambda_{\min}$ obtained for different system sizes: a critical point is present if this quantity tends to zero for increasing system sizes, while it is absent if if saturates to a finite value.
@@
Figure 2: (a) Illustration of the approach to the thermodynamic limit by expanding the size of a fractal pattern. At each step, the total linear size $L$ is multiplied by three, while the smallest resolved feature $l$ remains constant. (b–d): Minimal asymptotic decay rate $\lambda_{\min}$ as a function of the linear system size $L$ for detuning values $\Delta=1.0\gamma$, $\Delta=2.0\gamma$, and $\Delta=3.0\gamma$. Different markers correspond to different spatial geometries indicated in the legend. The grey dashed line indicates the numerical limit $1/t_{f}$, which corresponds to the frequency resolution of the simulations set by the finite evolution time $t_{f}$.
@@
Results at fixed detuning — We begin by examining the case $\Delta=\gamma$, for which in previous investigations the first-order phase transition was found to occur in 2D, and to be absent in 1D [15, 30, 20]. This is confirmed by our calculations of $\lambda_{\min}$, whose dependence on the system size is shown in Fig. 2(b), with green squares for a 2D system and orange bars for a 1D system. In fact, for the 2D case and in the limit of large systems, $\lambda_{\min}$ vanishes (with a precision set by the length $t_{f}$ of the time evolution). This critical slowing down is instead absent in the one-dimensional system, as $\lambda_{\min}$ saturates to a finite value.
@@
We extended this investigation to fractal dimensions. For example, for the Sierpinski carpet (second column of Fig. 1) the asymptotic decay rate has a minimum around the expected mean-field critical driving [1], but $\lambda_{\min}$ converges to a finite value for increasing system size, as shown in Fig. 2(b) with empty blue square markers. The system therefore exhibits slowing down, but it is not critical, so that, at least for this value of the detuning, the Sierpinski carpet of dimension $d_{H}\simeq 1.84$ is below the critical dimension. This reflects also on the steady-state average density, that does not display a discontinuous jump as a function of $I$, and has instead a smooth crossover from the low density to the high density phase [1]. We obtain a similar behavior for the Sierpinski cross (third column of Fig. 1), which has $d_{H}\simeq 1.5$ and whose convergence of $\lambda_{\min}$ to a finite value is shown with red cross markers in Fig. 2(b). It is interesting to comment that the shape the longest-lived mode associated to $\lambda_{\min}$ in a fractal system is non-trivial (see Figure in [1]) and will be the object of further studies.
@@
Figure 3: Minimal asymptotic decay rate $\lambda_{\min}$ versus Hausdorff dimension for a fixed detuning $\Delta=\gamma$. The patterns for the light blue circular markers are shown in [1]. Horizontal error bars show the uncertainty of the box-counting method, which is more significant only for patterns of fractal dimension smaller than one (Cantor sets).
@@
Besides the fact that it vanishes or not, the value of $\lambda_{\text{min}}$ also quantifies how far the system is from criticality. We found it to depend strongly on the system’s dimensionality, as shown in Fig. 3, where we plot $\lambda_{\text{min}}$ as a function of the Hausdorff dimension $d_{H}$, for a variety of fractal geometries. These values are the ones for the largest computalionally feasible sizes (ranging from $L=81l$ to $L=216l$, depending on the geometry), and are representative of the thermodynamic limit $\lim_{L\to\infty}\lambda_{\text{min}}$. In fact, the difference between $\lambda_{\min}$ for the two largest considered sizes gives us a vertical error bar that is always smaller than the marker size. Remarkably, $\lambda_{\min}$ exhibits a clear monotonous dependence on the Hausdorff dimension $d_{H}$. The slowing down continuously approaches a critical one while the dimension is increased, that in this case $\Delta=\gamma$ is only reached when the dimension is close to 2, i.e. the lower critical dimension for the first-order phase transition is $d_{L}(\Delta=\gamma)\approx 2$. This extends previous results [15, 30, 20] to fractional dimensions, but, as we show in the following, the lower critical dimension surprisingly decreases significantly from 2 as $\Delta$ increases.
@@
Figure 4: Minimal asymptotic decay rate as a function of detuning $\Delta$ for different spatial dimensions. All data points are obtained for a system size $L=81l$. The vertical error bars represent the difference in $\lambda_{\min}$ between this size and $L=27l$, providing an estimate of the finite-size effects and our proximity to the thermodynamic limit. Inset: Critical detuning $\Delta_{c}$ as a function of the Hausdorff dimension $d_{H}$. Error bars correspond to the numerical resolution. The gray dashed line shows the critical detuning predicted by the statistical mean-field approach (see main text), above which the phase transition is possible (shaded area).
@@
Changing the lower critical dimension — As shown in Figs. 2(c)-(d) and in Fig. 4, increasing the driving detuning $\Delta$ leads to a systematic decrease of $\lambda_{\min}$, and eventually to its vanishing, for all geometries. This implies that the system is critical and exhibits a first-order phase transition even for dimensions $d_{H}<2$, provided that the ratio $\Delta/\gamma$ is large enough. For a fixed dimension $d_{H}$, we define the critical detuning $\Delta_{c}$ as the value above which $\lambda_{\min}$ vanishes in the thermodynamical limit. For the 2D system (green squares), $\Delta_{c}\simeq\gamma$, close to the previously studied case. $\Delta_{c}$ is instead larger for lower-dimensional systems. This can be seen in Fig. 2(c), where, for large enough systems, $\lambda_{\min}$ vanishes not only for the 2D square, but also for the Sierpinski carpet, which for $\Delta=\gamma$ is not critical. Correspondingly, the smooth crossover in the average density for $\Delta=\gamma$, becomes a sharp discontinuity [1]. For $\Delta=2\gamma$ the lower-dimensional Sierpinski cross and the one-dimensional system are still non-critical, but both become for a higher $\Delta=3\gamma$, as shown in Fig. 2(d). These results show that it is possible to tune the lower critical dimension, or, in other words, to make the phase transition occur at any fixed dimension by increasing $\Delta$. The inset of Fig. 4, shows that the lower critical dimension decreases continuously as the detuning is increased. Interestingly, also the 1D case and fractals with dimension lower than one (e.g., Cantor sets) can be made critical by increasing $\Delta/\gamma$ enough. In fact, in the limit $d_{H}\to 0^{+}$, we can expect $\Delta_{c}\to+\infty$, since for a zero-dimensional (one-mode) nonlinear bosonic resonator it is known that the asymptotic decay rate vanishes only in the limit of infinite detuning [27, 10].
To complement our numerical results, we gain insight into the existence of a critical detuning by introducing an analytical statistical mean-field theory that incorporates the effect of kinetic energy contributions in the non-equilibrium Gross-Pitaevskii equation (1). For a fractal driving pattern, the system contains multiple interfaces between pumped and non-pumped regions, such that the kinetic energy contribution associated with these gradients becomes important. We therefore construct the statistical mean-field approximation by incorporating the kinetic contribution through the energy cost of a single interface, multiplied by the average number of interfaces within a region of size comparable to the smallest feature of the fractal pattern. This number can be related to the Hausdorff dimension through the average connectivity $\tilde{z}=2d_{H}$, while an analytical estimate of the kinetic contribution can be obtained from a solution of (1) with $g=0$. This derivation, detailed in the End Matter, yields an equation for the steady-state density with an effective, dimensionality-dependent detuning, which displays bistability when
$\Delta_{eff}=\Delta-\frac{4-2d_{H}}{4}\sqrt{\Delta^{2}+\frac{\gamma^{2}}{4}}>\frac{\sqrt{3}}{2}\gamma.$ (2)
This provides an estimate of the critical detuning, shown as the dashed gray line in the inset of Fig. 4. Since bistability is a necessary condition for the emergence of critical behavior, this prediction constitutes a lower bound for the critical detuning. It captures well the overall trend of the numerical data obtained from brute-force simulations of the full multimode dynamics, that lie slightly above this lower bound. This shows that, upon decreasing the dimensionality, the phase transition is suppressed at sufficiently low detuning, leading to a lower critical dimension that can be fractal.